OpenStudy (anonymous):

How do I find the particular solution for [e^(2x)]/(1+e^x) /i have a differential equation : y''' - 3y''+2y' = [e^(2x)]/(1+e^x) I understand how to find the complimentary solution but I am unsure of how to solve for the particular solution ( [e^(2x)]/(1+e^x) )

OpenStudy (freckles):

I heard of particular solutions for differential equations what is the particular solution meaning here?

OpenStudy (anonymous):

The particular solution used to determine a solution. Solution containing both complimentary and particular answers

OpenStudy (amistre64):

homogenous solution: de = 0 particular solution: de = f(x) when we know a homogenous solution, we can use a wronskian or variation of parameters to determine the particular solution

OpenStudy (amistre64):

not too sure how 'simple' that is to do here :)

OpenStudy (amistre64):

have you determined the homogenous solution?

OpenStudy (irishboy123):

.

OpenStudy (anonymous):

I found the complimentary (homogeneous) solution to be: x(t) = c1 e^x +c2 e^2x

OpenStudy (amistre64):

it y(t) that you are trying to determine y''' - 3y''+2y' = 0, let y' = u u'' - 3u'+2u = 0 r^2 -3r +2 = 0; when r=2 or 1 u = c0 e^(x) + c1 e^(2x) y' = c0 e^(x) + c1 e^(2x) y = c0 e^(x) + c1 e^(2x) + c3

OpenStudy (anonymous):

Thank you for the help!!! Would you be able to help me solve this for the non-homogeneous equation? I have no idea what to do :(

OpenStudy (amistre64):

sure, its a variation in the parameters .... let the undetermined constants be defined as undetermined functions of t then find y' y'' and y''' and insert them into the setup if you want an overview of what we are to do then i have a sort of tutorial: http://openstudy.com/users/amistre64#/updates/552fd761e4b0af026591bbbe

OpenStudy (anonymous):

Thank you so much for the tutorial!

OpenStudy (amistre64):

functions of x ... that is. your x(t) threw me :) we want to find y(x) which we did and can now use it to determine a particular solution. \[\hat y(x) = A e^x + B e^{2x} + C\] \[\hat y'(x) = A e^x + B \frac12e^{2x} +C+ (A' e^x + B' e^{2x} + C':=0)\] \[\hat y''(x) = A e^x + B \frac14e^{2x}+C+(A' e^x + B' \frac12e^{2x}+C':=0)\] \[\hat y'''(x) = A e^x + B \frac18e^{2x}+C+A' e^x + B' \frac14e^{2x}+C'\] now when we organize the data collected the tutrial i posted shows why we end up with a system of equations defined as: \[A' e^x + B' e^{2x} + C':=0\] \[A' e^x + B' \frac12e^{2x}+C':=0\] \[A' e^x + B' \frac14e^{2x}+C':= \frac{e^{2x}}{1+e^x}\]

OpenStudy (amistre64):

forgot to include that C e^(0x) derives to 0C and i posted all C instead ... but thats immaterial since it bears nothing on the solution.

OpenStudy (amistre64):

we could have done undetermined coefficients, but that tends to invoke an 'educated guess' that i simply do not have the tables for

OpenStudy (amistre64):

sigh .. seems i was thinking integration instead of derivative while trying to make some pretty coding..... 1/2 to 2 1/4 to 4

OpenStudy (anonymous):

We could also attempt to solve via reduction of order given that we know two linearly independent solutions.

OpenStudy (amistre64):

u = co e^x + c1 e^(2x) A' e^x + B' e^(2x) := 0 A' e^x + B' 2e^(2x) := e^(2x)/ (1+e^x) A' = -B' e^x -B' e^(2x) + B' 2e^(2x) := e^(2x)/ (1+e^x) B' [2e^(2x)-e^(2x)] := e^(2x)/ (1+e^x) B' e^(2x) := e^(2x)/ (1+e^x) ----------------------------- A' = -e^x/(1+e^x) B' = 1/(1+e^x) ---------------------------- A = -ln(1+e^x) ive go no good idea on B lol, wolf says: x - log(1+e^x) ------------------------------ u_p = -e^x ln(1+e^x) + (x - log(1+e^x)) e^(2x) y'_p = -e^x ln(1+e^x) + x e^(2x) - e^(2x) log(1+e^x) maybe .....

OpenStudy (amistre64):

working those integrals may or maynot be simple :)